pipe a takes 8 hours to empty out a pool. when added with pipe b it takes only 3 hourse. how long would pipe b take to empty out the pool?: pipe a takes 8 hours to empty out a pool. when added with pipe b it takes only 3 hourse. how long would pipe b take to empty out the pool?
Answer by ptaylor(1323)


Let x=amount of time it takes pipe B to empty out the pool
Then pipe B empties the pool at the rate of 1/x pool per hour
And pipe A empties the pool at the rate of 1/8 pool per hour
Together they empty the pool at the rate of (1/x + 1/8) pool per hour and we are told that, together, they empty the pool at the rate of 1/3 pool per hour
So our equation to solve is:
1/x + 1/8=1/3 multiply each term by 24x
24+3x=8x subtract 3x from each side
24+3x-3x=8x-3x
5x=24 divide each side by 5
x=4 4/5 hours------------------amount of time it takes pipe B to empty the pool
CK
(1/4.8) +1/8 = 1/3
5/24+3/24=1/3
8/24=1/3
1/3=1/3

Hope this helps---ptaylor